# Contrastive Learning from Synthetic Audio Doppelgängers

### v1

@@ Abstract
Learning robust audio representations currently demands extensive datasets of real-world sound recordings. By applying artificial transformations to these recordings, models can learn to recognize similarities despite subtle variations through techniques like contrastive learning. However, these transformations are only approximations of the true diversity found in real-world sounds, which are generated by complex interactions of physical processes, from vocal cord vibrations to the resonance of musical instruments. We propose a solution to both the data scale and transformation limitations, leveraging synthetic audio. By randomly perturbing the parameters of a sound synthesizer, we generate audio doppelgängers—synthetic positive pairs with causally manipulated variations in timbre, pitch, and temporal envelopes. These variations, difficult to achieve through transformations of existing audio, provide a rich source of contrastive information. Despite the shift to randomly generated synthetic data, our method produces strong representations, competitive with real data on standard audio classification benchmarks. Notably, our approach is lightweight, requires no data storage, and has only a single hyperparameter, which we extensively analyze. We offer this method111doppelgangers.media.mit.edu as a complement to existing strategies for contrastive learning in audio, using synthesized sounds to reduce the data burden on practitioners.
@@ 1 Introduction
“Noises have generally been thought of as indistinct, but that is not true.”
@@
The success of modern machine learning algorithms for tasks like audio understanding often hinges on both the quality and quantity of available data. Self-supervised learning methods, like contrastive learning, have even been able to leverage unlabeled data, enabling more human-like learning from patterns without needing explicit supervision. However, human perceptual processing is remarkably robust beyond this: for example, the human auditory system can easily recognize sounds across a wide range of variations, such as changes in pitch, timbre, or background noise. Moreover, humans can quickly learn to recognize novel sounds that they encounter in their environment. Replicating this ability to learn from a diverse array of sounds—or "noises," as we might call them—could significantly enhance the efficiency, scalability, and adaptability of machine learning models.
@@
Contrastive learning, which operates by recognizing similarities in the data among negative distractors, often relies on augmentations: transformations of input data that preserve content semantics. This method has been influential in audio representation learning, with specific implementations ranging from spectral masking to temporal jitter to cropping and other methods [1, 2, 3, 4, 5, 6, 7]. Data augmentations, though demonstrably useful, operate at the level of the observed data, not the underlying data-generating process as would be observed in real-world variation. They statistically alter data without directly manipulating the causal mechanisms that produced it, resulting in high correlation between augmented samples, as well as limited control and interpretability.
@@
In our work, we propose using a synthesizer to overcome this barrier, in addition to providing the scalability required for modern pretraining regimes through virtually unlimited data synthesis. A synthesizer can be understood as a system where parameters (relating to psychophysical attributes like pitch, timbre, and loudness) causally influence the generated sound. Modifying these parameters allows us to intervene in the data-generating process in a controllable way to generate positive pairs that vary in terms of their underlying synthesis parameters.
@@
We formulate an approach in which we randomly synthesize sounds, and then slightly perturb their parameters to generate positive pairs. We call these audio doppelgängers (examples in Contrastive Learning from Synthetic Audio Doppelgängers); they share a resemblance but are in fact distinct enough to learn from the variation between them. In a way, this approach uses an artificial data source effectively consisting of random synthetic noises but more “natural” augmentations akin to variation in similar sounds; as Pierre Schaeffer put it, noises are not indistinct. Through a comprehensive set of experiments, we show that models trained this way can yield strong performance on a wide range of downstream tasks, competitive with real audio datasets.
@@ 2.1 Learning from Synthetic Data
Synthetic data, artificially generated information rather than collected from real-world sources, has emerged as a valuable tool for learning across various domains [8, 9, 10, 11]. By addressing data scarcity, privacy concerns [12, 13], or removing biases [14, 15], synthetic data offers a promising avenue to complement real-world data and further drive progress in machine learning research.
Audio presents unique challenges due to the complexity of waveforms and temporal dependencies. Synthetic data has found applications in subareas like speech recognition [16, 17, 18, 19, 20] leveraging text-to-speech systems for detecting unspoken punctuation [21], recognizing low-resource languages [22], increasing acoustic diversity [23] or detecting out-of-vocabulary words [24]. However, non-speech audio domains can be highly diverse, requiring more complex approaches to data synthesis. In this domain, synthetic data has been used for specific tasks like timbre-text alignment [25] and vocoding [26]. The partially synthetic NSynth [27] dataset has also been used for pitch estimation and instrument classification. In our work, we tackle the general audio domain, proposing a synthetic data approach that can produce diverse sounds for general-purpose audio representation learning.
In computer vision, synthetic data is more popular and has been employed in different tasks to improve performance [28, 29, 30, 31, 32, 33, 34, 35]. While initially focused on using graphics engines to generate photorealistic scenes, recent work has investigated sampling synthetic data from deep generative models [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. However, these models aim to produce realistic images and still depend on real image datasets for training or synthesis. Thus, recent work has pushed away from realism, generating synthetic data such as fractals [49], or through other procedural noise models [50, 51] to use as training data for visual representation learners. In our work, we also abandon realism and leverage randomly generated synthetic sounds to learn audio representations for downstream tasks.
@@ 2.2 Contrastive Learning
A common strategy for learning from unlabeled data is contrastive learning. In this technique, we seek representations that are invariant to minor differences, i.e. they encode a space in which similar objects are closer together, and dissimilar objects are further. A classic strategy for this is to use data augmentations, transformations which noticeably alter a datapoint (for example, randomly cropping an image) without changing its essential content (e.g. what the image is of, such as a cat). These transformed versions then become a positive pair, while other examples (e.g. an image of a dog) become negatives. In audio, contrastive learning has been used extensively to produce high-quality representations for downstream tasks [5, 2, 52, 4, 53]. However, contrastive learning typically relies on augmentations with limited variation, by effecting already-existing material or sampling with offsets. Instead, our approach produces synthetic positive pairs that deviate through randomly perturbed parameters which causally influence acoustic properties of the output, such as the pitch of oscillators, or the temporal envelope of signals in a mixture. This allows us to mimic more naturalistic variation in auditory features without (or in addition to) applying post-hoc effects.
@@ 2.3 Sound Synthesis
The toolkit of sound synthesis has evolved to include a variety of hardware and software [54, 55, 56]. Synthesizers, abstractions of sound synthesis and processing methods often designed to act as musical instruments, are key to this: they expose control parameters that let sound designers guide them to produce desirable sounds for music, film, and many other applications. Accelerated synthesizers [57, 58] have recently allowed much faster-than-realtime sound generation, offering the ability to iteratively tweak parameters to reconstruct sounds [59, 60] and even match textual descriptions [61]. Such approaches highlight the practical utility of synthesizers: lightweight architectures controlled by a limited number of interpretable parameters are capable of producing a diverse array of sounds, often corresponding to well-known categories and concepts (e.g. the sound of waves can often be modeled with time-varying filtered noise). In our work, we leverage SynthAX [58], to rapidly produce diverse training data with controllable similarity between examples.
@@ 3.1 Data Generation
Our data generation pipeline uses virtual modular synthesizers implemented by SynthAX [58] in JAX. By default, we use the Voice synthesizer architecture [57], which has been shown to generate perceptually diverse sounds. This synthesizer has 78 parameters: a monophonic keyboard; two low-frequency oscillators and six ADSR temporal envelopes to generate control signals; sine and square-saw oscillators and a noise generator; and additional amplitude- and modulation-related components. In our experiments, we investigate two further architectures: VoiceFM has 130 parameters and includes a frequency modulation (FM) operator, and ParametricSynth has 2 sine and 2 square-saw oscillators, 1 sine FM and 1 square-saw FM operator, 340 in total. Varying the architecture allows us to investigate whether architectural complexity could affect the quality of representations learned. We generate 1-second sounds by default, for compatibility with most encoders (e.g. VGGish [62]). However, this practice can be extended to longer sounds.
@@
A key contribution of our work is synthesizing paired positive samples that sound alike, but are dissimilar due to their synthesis parameters and not only post-hoc effects (e.g. augmentations). This draws on the canonical definition from contrastive learning of positives that are sampled from the same latent class [63].
@@
In principle, $\delta$ controls the distance between the positive pairs and therefore the hardness of the contrastive learning task. Practically, we expect there to be a sweet-spot for $\delta$, considering prior work on mutual information and redundancy in contrastive learning problems [64, 65] as with very high $\delta$, the parameter vectors may become dominated by noise, resulting in difficulty effectively aligning their representations. Given this, we extensively study the effect of $\delta$ on downstream results.
@@ 3.2 Real Data
To compare to real audio data, we use sounds from VGGSound [66], a well-known dataset taken from YouTube videos (we only use audio). We use a random sample of 100,000 10-second files and select a random 1-second segment from each file at each iteration to augment. This allows us to fairly compare to our synthetic sounds by keeping duration constant, while still sampling from a variety of real sounds by randomizing the 1-second segments. Though VGGSound has labels included, we do not use them in training these models to keep the self-supervised constraint.
@@ 3.3 Preprocessing, Data Augmentations, and Audio Encoder
In our experiments, we use VGGish frontend representations [62]. We resample audio to 16kHz and obtain mel spectrograms with 64 mel bands and 96 time steps. We use a chain of effects as augmentations (implemented in torch-audiomentations [^2] https://github.com/asteroid-team/torch-audiomentations): a high-pass filter (cutoff frequency range 20–800Hz), a low-pass filter (1.2–8kHz), pitch shift (-2 to 2 semitones), time shift (-25% to 25%, rollover enabled), and finally reverberation for which we sample randomly from a set of impulse responses. All augmentations are applied with probability 0.5. We found that this yielded far stronger results than SpecAugment [67], and so we use this as a comparison point in all our experiments. More details on the augmentation are given in Appendix B. We also test temporal jitter, wherein different 1-second segments are sampled from within the same source clip and treated as positives [2, 3]. Our audio encoder is a ResNet18 [68], where we replace the initial layer with a 1-channel convolution to account for the effectively 1-channel spectrogram.
@@ 3.4 Contrastive Learning
We train for 200 epochs, generating (or sampling) 100,000 sounds per epoch, with a 90%-10% train-validation split. We use a batch size of 768 per GPU with two V100s. The training uses the alignment and uniformity objectives [69] used in prior work on learning with synthetic data [50]. We adopt the default parameters for these: $\textrm{unif}_{t}=2$, $\textrm{align}_{\alpha}=2$, and equal weights $\lambda_{1}=\lambda_{2}=1$ for both terms. Following this work, we use stochastic gradient descent for optimization, with a maximum learning rate of 0.72 (calculated as 0.12 ×$\times\dfrac{\textrm{total batch size}}{256}$) and weight decay $10^{-6}$. The learning rate follows a multi-step schedule with $\gamma=0.1$, and milestones at 77.5%, 85%, and 92.5% of the total learning epochs. Detailed steps are provided in Algorithm 1. Training with our synthetic data takes approx. 1-2 hours, as the data is generated on the fly in batches, whereas using on-disk datasets with effect chain augmentations can extend training time up to 6-8+ hours.
Algorithm 1 Our contrastive learning procedure with audio doppelgängers. In the training loop, we drop the batch index $i$ for simplicity.
@@
0:  $\ell_{\textrm{unif}}(\textbf{X}\in[0,1]^{k\times m_{M}})\leftarrow\log\left[\dfrac{1}{k^{2}}\exp\left({-t\sum\limits_{j=1}^{k}\sum\limits_{l=1}^{k}\|\mathbf{X}[j]-\mathbf{X}[l]\|_{2}^{2}}\right)\right]$ where $t=2$

### v2

@@ Abstract
Learning robust audio representations currently demands extensive datasets of real-world sound recordings. By applying artificial transformations to these recordings, models can learn to recognize similarities despite subtle variations through techniques like contrastive learning. However, these transformations are only approximations of the true diversity found in real-world sounds, which are generated by complex interactions of physical processes, from vocal cord vibrations to the resonance of musical instruments. We propose a solution to both the data scale and transformation limitations, leveraging synthetic audio. By randomly perturbing the parameters of a sound synthesizer, we generate audio doppelgängers—synthetic positive pairs with causally manipulated variations in timbre, pitch, and temporal envelopes. These variations, difficult to achieve through augmentations of existing audio, provide a rich source of contrastive information. Despite the shift to randomly generated synthetic data, our method produces strong representations, outperforming real data on several standard audio classification tasks. Notably, our approach is lightweight, requires no data storage, and has only a single hyperparameter, which we extensively analyze. We offer this method as a complement to existing strategies for contrastive learning in audio, using synthesized sounds to reduce the data burden on practitioners.
@@ 1 Introduction
“Noises have generally been thought of as indistinct, but this is not true.”
@@
The success of modern machine learning algorithms for tasks like audio understanding often hinges on both the quality and quantity of available data. Self-supervised learning methods, like contrastive learning, have even been able to leverage unlabeled data, enabling more human-like learning from patterns without needing explicit supervision. However, human perceptual processing is remarkably robust beyond this: for example, the human auditory system can easily recognize sounds across a wide range of variations, such as changes in pitch, timbre, or background noise. Moreover, humans can quickly learn to recognize novel sounds that they encounter in their environment. Replicating this ability to learn from a diverse array of sounds—or ”noises,” as we might call them—could significantly enhance the efficiency, scalability, and adaptability of machine learning models.
@@
Contrastive learning, which operates by recognizing similarities in the data among negative distractors, often relies on augmentations: transformations of input data that preserve content semantics. This method has been influential in audio representation learning, with specific implementations ranging from spectral masking to temporal jitter to cropping and other methods (Huang et al., 2022; Saeed et al., 2021; Spijkervet & Burgoyne, 2021; Wang & Oord, 2021; Al-Tahan & Mohsenzadeh, 2021; Niizumi et al., 2021; Manocha et al., 2021). Data augmentations, though demonstrably useful, operate at the level of the observed data, not the underlying data-generating process as would be observed in real-world variation. They statistically alter data without directly manipulating the causal mechanisms that produced it, resulting in high correlation between augmented samples, as well as limited control and interpretability.
@@
In our work, we propose a different strategy: using a synthesizer to overcome this barrier, in addition to providing the scalability required for modern pretraining regimes through virtually unlimited data synthesis. A synthesizer can be understood as a system where parameters (relating to psychophysical attributes like pitch, timbre, and loudness) causally influence the generated sound. Modifying these parameters allows us to intervene in the data-generating process in a controllable way to generate positive pairs that vary in terms of their underlying synthesis parameters. Unlike traditional data augmentation techniques, our method generates entirely synthetic audio data from scratch. This approach allows us to control the underlying data-generating process directly, offering a perspective distinct from augmentation of real data.
@@
We formulate an approach in which we randomly synthesize sounds, and then slightly perturb their parameters to generate positive pairs. We call these audio doppelgängers (examples in Contrastive Learning from Synthetic Audio Doppelgängers); they share a resemblance but are in fact distinct enough to learn from the variation between them. In a way, this approach uses an artificial data source effectively consisting of random synthetic noises but more “natural” differences akin to variation in similar sounds; as Pierre Schaeffer put it, noises are not indistinct. Through a comprehensive set of experiments, we show that models trained this way can yield strong performance on a wide range of downstream tasks, competitive with real audio.
@@ 2.1 Learning from Synthetic Data
Synthetic data, artificially generated information rather than collected from real-world sources, has emerged as a valuable tool for learning across various domains (Liu et al., 2023; Silver et al., 2017; Kumar et al., 2020; Meng et al., 2022). By addressing data scarcity, privacy concerns (Tucker et al., 2020; DuMont Schütte et al., 2021), or removing biases (Tan et al., 2020; Ramaswamy et al., 2021), synthetic data offers a promising avenue to complement scarce Longpre et al. (2024) real-world data and further drive progress in machine learning research.
Audio presents unique challenges due to the complexity of waveforms and temporal dependencies. Synthetic data has found applications in subareas like speech recognition (Rosenberg et al., 2019; Rossenbach et al., 2020; Laptev et al., 2020; Fazel et al., 2021; Hu et al., 2022; Gao et al., ) leveraging text-to-speech systems for detecting unspoken punctuation (Soboleva et al., 2021), recognizing low-resource languages (Bartelds et al., 2023), increasing acoustic diversity (Chen et al., 2020b) or detecting out-of-vocabulary words (Zheng et al., 2021). However, non-speech audio domains can be highly diverse, requiring more complex approaches to data synthesis. In this domain, synthetic data has been used for specific tasks like timbre-text alignment (Jonason & Sturm, 2022) and vocoding (Wang et al., 2023). The partially synthetic NSynth (Engel et al., 2017) dataset has also been used for pitch estimation and instrument classification. In our work, we tackle the general audio domain, proposing a synthetic data approach that can produce diverse sounds for general-purpose audio representation learning.
In computer vision, synthetic data is more popular and has been employed in different tasks to improve performance (Chen et al., 2019; Ros et al., 2016; Varol et al., 2017; Ionescu et al., 2013; Shakhnarovich et al., 2003; Mayer et al., 2016; Dosovitskiy et al., 2015; Ren & Lee, 2018). While initially focused on using graphics engines to generate photorealistic scenes, recent work has investigated sampling synthetic data from deep generative models (Besnier et al., 2020; Ravuri & Vinyals, 2019; Jahanian et al., 2021; Zhang et al., 2021; Tritrong et al., 2021; Li et al., 2021; Shrivastava et al., 2017; Hoffman et al., 2018; Tian et al., 2024; 2023; Trabucco et al., 2023; Yang et al., 2021; Jeong & Shin, 2021). However, these models aim to produce realistic images and still depend on real image datasets for training or synthesis. Thus, recent work has pushed away from realism, generating synthetic data such as fractals (Kataoka et al., 2020), or through other procedural noise models (Baradad Jurjo et al., 2021; Baradad et al., 2022) to use as training data for visual representation learners. In our work, we also abandon realism and leverage randomly generated synthetic sounds to learn audio representations for downstream tasks.
@@ 2.2 Contrastive Learning
A common strategy for learning from unlabeled data is contrastive learning. In this technique, we seek representations that are invariant to minor differences, i.e. they encode a space in which similar objects are closer together, and dissimilar objects are further. A classic strategy for this is to use data augmentations, transformations which noticeably alter a datapoint (for example, randomly cropping an image) without changing its essential content (e.g. what the image is of, such as a cat). These transformed versions then become a positive pair, while other examples (e.g. an image of a dog) become negatives. In audio, contrastive learning has been used extensively to produce high-quality representations for downstream tasks (Al-Tahan & Mohsenzadeh, 2021; Saeed et al., 2021; Ravanelli & Bengio, 2019; Wang & Oord, 2021; Fonseca et al., 2021b). Our approach differs fundamentally from data augmentation strategies commonly used in contrastive learning. Instead of applying transformations to existing audio samples, we generate synthetic audio pairs by perturbing synthesizer parameters, creating positive pairs with causal variations that are difficult to achieve through augmentations. This represents a novel application of synthetic data in the context of general-purpose audio representation learning.
@@ 2.3 Sound Synthesis
The toolkit of sound synthesis has evolved to include a variety of hardware and software (Mathews et al., 1969; Pinch & Trocco, 2004; Théberge, 1997). Synthesizers, abstractions of sound synthesis and processing methods often designed to act as musical instruments, are key to this: they expose control parameters that let sound designers guide them to produce desirable sounds for music, film, and many other applications. Accelerated synthesizers (Turian et al., 2021; Cherep & Singh, 2023) have recently allowed much faster-than-realtime sound generation, offering the ability to iteratively tweak parameters to reconstruct sounds (Hagiwara et al., 2022; Shier, 2021) and even match textual descriptions (Cherep et al., 2024). Such approaches highlight the practical utility of synthesizers: lightweight architectures controlled by a limited number of interpretable parameters are capable of producing a diverse array of sounds, often corresponding to well-known categories and concepts (e.g. the sound of waves can often be modeled with time-varying filtered noise). In our work, we leverage SynthAX (Cherep & Singh, 2023), to rapidly produce diverse training data with controllable similarity between examples.
@@ 3.1 Data Generation
Our data generation pipeline uses virtual modular synthesizers implemented by SynthAX (Cherep & Singh, 2023) in JAX. By default, we use the Voice synthesizer architecture (Turian et al., 2021), which can generate perceptually diverse sounds. Our synthesizer consists of several common modules (with parameter-counts in parenthesis):
  • Keyboard (2x): Controls the sound’s fundamental frequency ($f_{0}$) and duration.
  • Low-Frequency Oscillators (LFOs; 8x each): Two LFOs modulate various aspects of the sound, each with parameters for frequency, modulation depth, initial phase, and amplitude weights across waveforms.
  • ADSR Envelopes (5x each): Six envelopes shape the amplitude and modulation signals, each defined by attack, decay, sustain, release, and curvature ($\alpha$).
  • Voltage-Controlled Oscillators (VCOs): Includes a sine VCO (3x) with tuning, modulation depth, initial phase, and a square-saw VCO (4x) adding waveform shape.
  • Noise Generator: Provides broadband noise without additional parameters.
  • Modulations (20x): Weight matrix controlling how modulation sources affect destinations.
  • Audio Mixer (3x): Combines outputs of oscillators and noise generator.
In total, the Voice synthesizer has 78 parameters. Perturbing these parameters allows us to generate a wide variety of sounds with controlled variations, such as slightly lower or higher pitch, a slightly longer onset or release, or a little more or less noise. In our experiments, we investigate two further architectures: VoiceFM has 130 parameters and includes a frequency modulation (FM) operator, and ParametricSynth has 2 sine and 2 square-saw oscillators, 1 sine FM and 1 square-saw FM operator, 340 in total. Varying the architecture allows us to investigate whether architectural complexity could affect the quality of representations learned. We generate 1-second sounds by default, for compatibility with most encoders (e.g. VGGish (Hershey et al., 2017)). However, this practice can be extended to longer sounds.
@@
A key contribution of our work is synthesizing paired positive samples that sound alike, but are dissimilar due to their synthesis parameters and not only post-hoc effects (e.g. augmentations). This draws on the canonical definition from contrastive learning of positives that are sampled from the same latent class (Saunshi et al., 2019).
@@
In principle, $\delta$ controls the distance between the positive pairs and therefore the hardness of the contrastive learning task. Practically, we expect there to be a sweet-spot for $\delta$, considering prior work on mutual information and redundancy in contrastive learning problems (Tian et al., 2020; Tosh et al., 2021) as with very high $\delta$, the parameter vectors may become dominated by noise, resulting in difficulty effectively aligning their representations. Given this, we extensively study the effect of $\delta$ on downstream results.
@@ 3.2 Real Data
To compare to real audio data, we use sounds from VGGSound (Chen et al., 2020a), a well-known dataset taken from YouTube videos (we only use audio). We use a random sample of 100,000 10-second files and select a random 1-second segment from each file at each iteration to augment. This allows us to fairly compare to our synthetic sounds by keeping duration constant, while still sampling from a variety of real sounds by randomizing the 1-second segments. Though VGGSound has labels included, we do not use them in training these models to keep the self-supervised constraint. Note that VGGSound is currently one of the largest publicly released audio datasets for pretraining, unlike AudioSet (which only releases URLs, not the content itself).
@@ 3.3 Preprocessing, Data Augmentations, and Audio Encoder
In our experiments, we use VGGish frontend representations (Hershey et al., 2017). We resample audio to 16kHz and obtain mel spectrograms with 64 mel bands and 96 time steps. We use a chain of effects as augmentations (implemented in torch-audiomentations [^1] https://github.com/asteroid-team/torch-audiomentations): a high-pass filter (cutoff frequency range 20–800Hz), a low-pass filter (1.2–8kHz), pitch shift (-2 to 2 semitones), time shift (-25% to 25%, rollover enabled), and finally reverberation for which we sample randomly from a set of impulse responses. All augmentations are applied with probability 0.5. We found that this yielded far stronger results than SpecAugment (Park et al., 2019), and so we use this as a comparison point in all our experiments. More details on the augmentation are given in Appendix B. We also test temporal jitter, wherein different 1-second segments are sampled from within the same source clip and treated as positives (Saeed et al., 2021; Spijkervet & Burgoyne, 2021). Our audio encoder is a ResNet18 (He et al., 2016), where we replace the initial layer with a 1-channel convolution to account for the effectively 1-channel spectrogram.
@@ 3.4 Contrastive Learning
We train for 200 epochs, generating (or sampling) 100,000 sounds per epoch, with a 90%-10% train-validation split. We use a batch size of 768 per GPU with two V100s. The training uses the alignment and uniformity objectives (Wang & Isola, 2020) used in prior work on learning with synthetic data (Baradad Jurjo et al., 2021). We adopt the default parameters for these: $\textrm{unif}_{t}=2$, $\textrm{align}_{\alpha}=2$, and equal weights $\lambda_{1}=\lambda_{2}=1$ for both terms. Following this work, we use stochastic gradient descent for optimization, with a maximum learning rate of 0.72 (calculated as 0.12 ×$\times\dfrac{\textrm{total batch size}}{256}$) and weight decay $10^{-6}$. The learning rate follows a multi-step schedule with $\gamma=0.1$, and milestones at 77.5%, 85%, and 92.5% of the total learning epochs. Detailed steps are provided in Algorithm 1. Training with our synthetic data takes approx. 1-2 hours, as the data is generated on the fly in batches, whereas using on-disk datasets with effect chain augmentations can extend training time up to 6-8+ hours.
Algorithm 1 Our contrastive learning procedure with audio doppelgängers. In the training loop, we drop the batch index $i$ for simplicity. We also show the pairwise distance in $\ell_{\textrm{nuif}}$, though the implementation (via torch.pdist) uses a condensed representation.
@@
0:  $\ell_{\textrm{unif}}(\textbf{X}\in\mathbb{R}^{k\times m_{M}})\leftarrow\log\left[\dfrac{1}{k^{2}}\sum\limits_{j=1}^{k}\sum\limits_{l=1}^{k}\exp\left({-t\|\mathbf{X}[j]-\mathbf{X}[l]\|_{2}^{2}}\right)\right]$ where $t=2$